Solve for z: z2-8z=9

Found 2 solutions by Cromlix, stanbon:
Answer by Cromlix(4381)

z^2 - 8z - 9 )= 0
(z - 9)(z + 1) = 0
z = 9
OR
z = -1
Hope this helps.
:-)

Answer by stanbon(75887)

z2-8z=9
------
z^2 - 8z - 9 = 0
Factor:
----
(z-9)(z+1) = 0
---
Ans: z = 9 or z = -1
========================
Cheers,
Stan H.
===============